Evlauate:
P(5,0)
Evlauate the expression
C(8,3)
Evaluate:
-5!
Evaluate:
5!-2!

Answer by stanbon(75887)

Evlauate:
P(5,0)
= 5!/(5-0)! = 1
---------------------------
Evlauate the expression
C(8,3)
= 8!/[(8-3)!*3!] = (8*7*6)(1*2*3)= 56
---------------------------------------------
Evaluate:
-5! = -120

Evaluate:
5!-2! = 120-2 = 118
========================
Cheers,
Stan H.
==============